OpenStudy (anonymous):

The fourth term of an arithmetic sequence is 141, and the seventh term is 132. The first term is _____.

OpenStudy (anonymous):

\[a, a+d,a+2d,a+3d,a+4d,...\] you got \(a+3d=141\) and \(a+6d=132\)

OpenStudy (anonymous):

that means \(132-141=(a+6d)-(a+3d)=3d=-9\) and so \(d=-3\)

OpenStudy (anonymous):

okay so then d=4?

OpenStudy (anonymous):

actually \(d=-3\)

OpenStudy (anonymous):

umm please explain!!

OpenStudy (anonymous):

k lets go slow

OpenStudy (anonymous):

please and thanks

OpenStudy (anonymous):

actually before we even start, since \(132<141\) is it clear that the terms are getting smaller?

OpenStudy (anonymous):

in other words, "\(d\)" the "common difference" must be negative right?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

if we call the first term \(a\) then the second term is \(a+d\) for some \(d\) and the third term is \(a+2d\), the fourth term is \(a+3d\) the fifth term is \(a+4d\) etc

OpenStudy (anonymous):

in other words, you keep adding \(d\) to each term to get the next term, with the sophistication that you might be "adding" a negative number

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The fourth term of an arithmetic sequence is 141 tells you that \(a+3d=141\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you see that it is the fourth term, so it is \(a+3d\) not \(a+4d\)

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

and the seventh term is 132 means \[a+6d=132\]

OpenStudy (anonymous):

from these two pieced of information we can solve for \(d\) and then solve for \(a\)

OpenStudy (anonymous):

*pieces

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

a bit of algebra shows that \[a+6d-(a+3d)=3d\] right?

OpenStudy (anonymous):

ohh okay i get it..

OpenStudy (anonymous):

so we see that \[3d=132-141=-9\]

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

and so since it is 7-4= 3 then -9/3 would be -3 right? giving us the difference

OpenStudy (anonymous):

yeah \(-3\) is the difference

OpenStudy (anonymous):

what you said

OpenStudy (anonymous):

ohhh okay!!! i get it!! :D

OpenStudy (anonymous):

you are still not done though right?

OpenStudy (anonymous):

your question asked "The first term is _____"

OpenStudy (anonymous):

hmm well pluggin in the difference and then your equation... a4=141 a3=141+3=144 a2=144+3=147 a1=147+3=150 so then the first term would be 150 right? i just did it backwards

OpenStudy (anonymous):

yeah i guess so i would have said \(a+3\times (-3)=141\)or \[a-9=141\] making \(a=150\) your method means you understand what is going on, which is good but you certainly wouldn't want to use that if you had say \(a_{75}\) and wanted \(a_1\)

OpenStudy (anonymous):

:D yay thank you!!! :D ohh okay... ill keep in mind that equation!! thank you soo much!

OpenStudy (anonymous):

yw